OpenStudy (anonymous):

ind_int xcos(5x) dx u=x du=dx dv=cos(5x) dx v=u-sub {u=5x du=5 v=ind_int cos(5x) =sin(u) du =sin(5x)*5 (1/5)v=sin(5x) My algebra is retricebackwards. Can someone please help me clear this up so I can better understand what I'm doing. Especially with notation and whatnot. thanks.

mathslover (mathslover):

can u write this with equation forum

OpenStudy (anonymous):

\[\int\limits_{}^{}xcos(5x)dx\]

OpenStudy (anonymous):

My v should equal (1/5)sin(5x), but I'm unclear on the notations and it makes the algebra confusing.

OpenStudy (anonymous):

v=\[\int\limits_{}^{}\cos(5x)\]

OpenStudy (anonymous):

v=\[\int\limits_{}^{}\cos(5x)dx\]

OpenStudy (anonymous):

5x=u d(5x)/dx=du/dx 5=du/dx du=5dx

OpenStudy (anonymous):

\[\int xcos(5x)dx\]\[=\frac{1}{5}\int xd(sin5x)\]Integration by parts\[=\frac{1}{5}[xsin5x-\int sin5xd(x)]\]\[=\frac{1}{5}[xsin5x-(-\frac{1}{5}cos5x)]+C\]\[=\frac{1}{5}xsin5x+\frac{1}{25}cos5x+C\]Not sure if it is correct

OpenStudy (anonymous):

therefore cos5xdx=(1/5)(5cos5x)dx=(1/5)(cos5x) 5dx=(1/5cosu)du

OpenStudy (anonymous):

That's the right ans, but I have problems with the notation in my u-sub. There's u's, du's, v's and dx's, and when I'm solving it out algebraically the meanings of the notations get lost to me. It's hard to convey across a forum.

OpenStudy (anonymous):

if you want i could send you some scanned texts to your email

OpenStudy (anonymous):

Do yo understand the first step in my answer?

OpenStudy (anonymous):

@Chumku I think you can upload the scanned text here via the ''attach file'' function.

OpenStudy (anonymous):

ya sure

OpenStudy (anonymous):

but i will be able to attach the file around after three hours or so